# Towards Comprehensive and Efficient Post Safety Alignment of Large Language Models via Safety Patching

### v1

@@ Appendix B SNIP Score for Identifying Important Neurons in Patches
Following Wei et al. [2024], we calculate the individual score for each data point and aggregate over the all examples. Intuitively, $S(W)$ measures the importance of each parameter entry in influencing the model’s behavior on the harmful dataset $D_{h}$ with different gradient optimization techniques. A small $S(W_{ij})$ indicates that setting $W_{ij}$ to zero has a negligible impact on the harmful data points $x$, allowing us to attribute specific characteristics of the model to the weights with large $S(W)_{ij}$.
@@
Table 4: There are six response categories. Responses in categories 0 to 4 are generally harmless, while those in category 5 are harmful [Wang et al., 2023].
Appendix C Automatic Evaluation of Harmful Rate
As pointed out by Wang et al. [2023], a key limitation of GPT-4-based evaluation is data privacy, as the model cannot be deployed locally. To address this problem, they introduce PLM-based evaluators. To be more specific, a Longformer-based [Beltagy et al., 2020] classifier is fine-tuned over human annotations for each input-output harmful data to perform harmful response detection. Each response to be evaluated will be classified into six categories as shown in Table 4, with only Lable 5 being considered as a harmful response.
Appendix D Details for Baselines
D.1 Baselines for Post Safety Alignment
We evaluate SafePatching against the following baseline methods.
Unlearning-based:
@@ D.1 Baselines for Post Safety Alignment
  • Gradient Ascent (GA) [Yao et al., 2023]: Gradient ascent is performed on harmful dataset during the training process. Furthermore, it also applies a forward KL-divergence with the original backbone model on extra general datafor utility keeping.
@@
  • GA + Mismatch [Yao et al., 2023]: Built upon GA approach, GA + Mismatch adds random responses from normal dataset for harmful inputs. And another gradient descent optimization is added accordingly to each training step. The forward KL-divergence is also leveraged on extra general data for the preservation of utility.
  • Task Vector [Zhang et al., 2023b]: The original backbone model is fine-tuned on the harmful dataset with gradient descent. Then we extract the harmful parameters (difference between the fine-tuned model and the original backbone) and perform negation.
@@
  • Negative Preference Optimization (NPO) [Zhang et al., 2024a]: A simple preference-optimization-inspired method that could efficiently and effectively unlearn a target dataset with negative samples. The forward KL-divergence is also applied on extra general data.
Decoding-based:
  • Self-Contrastive Decoding (Self-CD) [Shi et al., 2024]: A decoding-based method for over-safety mitigation. It collects the model’s responses to the same question by emphasizing within the prompts whether safety is taken into account. Then it mitigates over-safety by contrasting the output distributions of these answers.
@@
  • ROSE [Zhong et al., 2024]: A simple PSA method to directly boost the safety of existing aligned LLMs. The principle of it is to improve the probability of desired safe output via suppressing the undesired output induced by the carefully-designed reverse prompts.
@@
  • SafeDecoding [Xu et al., 2024]: It is designed to mitigate jailbreak attacks by identifying safety disclaimers and amplifying their token probabilities, while simultaneously attenuating the probabilities of token sequences that are aligned with the objectives of jailbreak attacks.
D.2 Baselines for Model Merging
@@
To further demonstrate the effectiveness of our SafePatching in mitigating the conflict between patches for safety and over-safety, while preserving the utility of the backbone model, we compare it with the current state-of-the-art model merging methods.
@@
  • Average Merging: It combines the parameters of multiple models by averaging them to create a merged model.
@@
  • Task Arithmetic [Ilharco et al., 2022a]: It uses a scaling factor to balance the contributions between the backbone model and the models being merged.
@@
  • Fisher Merging [Matena and Raffel, 2022]: First, it estimates the significance of the parameters by calculating the Fisher information matrix, and then it merges the parameters according to their importance.
@@
  • TIES-Merging [Yadav et al., 2023]: It seeks to resolve parameter conflicts during model merging. Initially, it trims parameters with lower magnitudes and then addresses sign disagreements. Finally, parameters with consistent signs are merged.
Appendix E Implementation Details
@@
Our experiments are implemented with PyTorch [Paszke et al., 2019] and Transformer library [Wolf et al., 2020]. LLaMA-2-Chat-7B and LLaMA-2-Chat-13B are performed on 4 NVIDIA Tesla A100 using DeepSpeed [Rasley et al., 2020] repository. And the evaluation for utility benchmarks is performed with lm-evaluation-harness [Gao et al., 2023].
Experimental Setting for Vanilla Harmful Prompt
The harmful training data $D_{h}$ consists a subset of the original training set from Beavertails [Ji et al., 2023], where up to 100 instances from the 14 safety categories are randomly sampled, leading to 1,385 training instances in total. And the testing set is also randomly sampled from the test split of Beavertails, with up to 100 instances per category (1,207 in total). And the testing instances in AdvBench is 520. We train SafePatching for 1 epoch with gradient ascent and 3 epochs with gradient decent, with the total batchsize of 32 and the constant learning rate of 5e-6 and 5e-5, respectively. The overall retention rate, most important neuron rate of safety and over-safety patches are set to 30%, 3% and 2%, respectively. As for the hyper-parameter $\alpha$ and $\beta$, they are set to 1 and 0.6.
@@
Further, we carefully evaluate the official implementations of all baselines, in order to make the comparison as fair as possible. We strictly follow the hyper-parameter settings in their original code, where maximum training step of gradient ascent is set to 1,000. Following [Yao et al., 2023], TruthfulQA [Lin et al., 2022] is adopted for the general training data $D_{g}$. If this could not reach the expected performance, we carry out the hyper-parameter search of the learning rate and batchsize for them. All the methods are evaluated for 3 random runs.
Experimental Setting for Jailbreak Attack
Here the harmful training data $D_{h}$ is formed by those harmful questions that are successfully jailbroken by ReNeLLM [Ding et al., 2023]. This is a simulation of SafePatching for “post-hoc” remediation of the jailbreak attack. Specifically, we collect 1,381 jailbroken data in total. We train SafePatching for 1 epoch with gradient ascent and 3 epochs with gradient decent, with the total batchsize of 32 and the constant learning rate of 5e-6 and 5e-5, respectively. The overall retention rate, most important neuron rate of safe and over-safety patches are set to 30%, 3% and 2%, respectively. As for the hyper-parameter $\alpha$ and $\beta$ in Equation (8), they are set to 1 and 1. The evaluation is performed on the AdvBench under the attack of ReNeLLM.
Experimental Setting for Continual Safety Alignment
We select three safety categories from Beavertails to simulate user safety alignment requirements at three consecutive time steps. Specifically, these three safety categories are “drug_abuse,weapons,banned_substance”, “animal_abuse”, and “misinformation_regarding_ethics,laws_and_safety”, each with 1,000 harmful data points for training and 100 for testing. At each time step, we train SafePatching for 1 epoch with gradient ascent and 3 epochs with gradient decent, with the total batchsize of 32 and the constant learning rate of 1e-5 and 5e-5, respectively. The overall retention rate, most important neuron rate of safe and over-safety patches are set to 30%, 3% and 2%, respectively. As for the hyper-parameter $\alpha$ and $\beta$ in Equation (8), they are set to 1 and 1. In Table 3, we report the average harmful rate on Beavertails, harmful refusal rate on XSTest and performance score on MT-Bench over the three time steps. And in Figure 5, the fine-grained performance changes over the continual alignment process are depicted.
Experimental Setting for Human Evaluation on Over-Safety
@@
Following Röttger et al. [2023], the LLM backbone would generate three main types of responses on the XSTest and OKTest: (1) Full Compliance, which is the preferred outcome for safe prompts. This includes any direct attempt to answer the question, regardless of accuracy or relevance. (2) Full Refusal, which is an undesirable outcome for safe prompts. In this case, the model explicitly states that it will not answer the question and may offer alternative suggestions instead. (3) Partial Refusal, which combines elements of both refusal and compliance. This includes responses that initially refuse to answer but then go on to address the question. And we only count responses fall into the Full Compliance as the refusal rate to measure the performance of over-safety. Besides, the participants in our human evaluation are volunteered transparently informed of our research intent, with reasonable wages paid.
Table 5: The detailed results on the utility benchmarks with LLaMA-2-Chat-7B backbone.
ARC-C HellaSwag MMLU WinoGrande BBH GSM8K MATH
Original 44.20 57.79 46.38 66.46 40.42 22.29 4.88
GA 42.24 55.48 46.30 66.22 38.69 21.68 4.69
GA + Mismatch 40.10 55.75 45.86 66.46 39.92 22.14 4.78
Task Vector 35.58 50.53 44.74 62.43 37.37 21.23 4.08
NPO 41.30 55.68 46.15 64.40 39.35 21.74 4.82
Average Merging 45.56 58.17 45.41 65.90 40.07 21.38 4.68
Task Arithmetic 45.31 58.13 45.83 66.06 40.44 22.14 5.32
Fisher Merging 45.99 57.95 44.56 66.14 38.37 12.51 3.94
TIES-Merging 45.05 58.16 45.27 65.35 40.33 21.99 5.38
SafePatching 44.12 57.78 45.11 66.82 39.35 21.99 4.76
Table 6: The detailed results on the utility benchmarks with LLaMA-2-Chat-13B backbone.
ARC-C HellaSwag MMLU WinoGrande BBH GSM8K MATH
Original 46.08 60.72 53.15 70.96 47.78 36.54 5.24
GA 46.08 54.36 52.86 69.38 46.6 36.09 5.50
GA + Mismatch 43.94 57.56 53.03 70.09 46.51 35.86 5.36
Task Vector 43.94 58.98 52.28 69.77 47.1 31.61 5.46
NPO 44.97 60.94 52.81 69.85 47.04 34.34 5.70
SafePatching 44.62 60.09 53.06 70.32 47.49 32.6 5.70
Table 7: The overall results on the safety, over-safety and utility benchmarks with LLaMA-2-Chat-13B backbone. The evaluation metrics for safety and over-safety is harmful rate and refusal rate, respectively. We report the average results on all the utility dataset except for MT-bench (1 - 10).
@@
AdvBench Beavertails XSTest OKTest AVG. MT-Bench
Original 23.08 12.34 8.40 2.67 45.78 6.47
GA 1.35 2.90 34.40 17.67 44.41 5.32
GA + Mismatch 14.42 11.10 12.40 5 44.62 6.49
Task Vector 2.50 1.07 62 55 44.16 5.95
NPO 29.40 16.60 9.40 8.67 45.09 5.50
SafeDecoding - - - - - -
Self-CD 0 0.02 25.60 31.33 - 4.63
ROSE 0 0.02 26.80 36.33 - 4.44
SafePatching 4.32 11.02 6.80 1 44.84 6.63
Table 8: Types of training data required by different methods.
@@
Table 9: ATGR of different methods to manifest the inference cost.
@@
Appendix F Additional Experimental Results
F.1 Detailed Results on Utility Benchmarks
We present the experimental results of baseline methods on each utility benchmark with different scales of LLaMA-2 backbone in Table 5 and 6. It can be seen that our SafePatching maintains good generalizability across different model scales. We do not report the results of these decoding-based methods on the utility benchmark due to their enormous inference costs, which make the entire testing process highly inefficient. The loss of the utility of the backbone model can be reflected in the performance degradation on MT-bench.
F.2 Results on LLaMA-2-Chat-13B backbone
The experimental results of the current PSA baseline methods based on unlearning and decoding on the LLaMA-2-Chat-13B backbone model are shown in Table 7. The results demonstrate that our SafePatching is still capable of achieving the three goals of enhancement of safety, mitigation of over-sensitivity, and maintenance of utility. Similarly, SafePatching continues to improve the model’s multi-turn instruction-following ability on the MT-bench, enhancing the quality of the model’s responses to user inputs by reducing its over-sensitivity to safety concerns. We hope that our method can provide inspiration for future work in achieving comprehensive PSA. We do not report the results of SafeDecoding on the LLaMA-2-Chat-13B backbone because it requires using a portion of safe data to train a safety-enhanced expert model. However, this portion of data is not publicly available, so we are unable to reproduce the results.
F.3 Comparison of Training and Inference Costs

### v2

@@ Appendix B SNIP Score for Identifying Important Neurons in Patches
Following Wei et al. (2024), we calculate the individual score for each data point and aggregate over the all examples. Intuitively, $S(W)$ measures the importance of each parameter entry in influencing the model’s behavior on the harmful dataset $D_{h}$ with different gradient optimization techniques. A small $S(W_{ij})$ indicates that setting $W_{ij}$ to zero has a negligible impact on the harmful data points $x$, allowing us to attribute specific characteristics of the model to the weights with large $S(W)_{ij}$.
Table 4: There are six response categories. Responses in categories 0 to 4 are generally harmless, while those in category 5 are harmful (Wang et al., 2023).
@@
Appendix C Details for Baselines
C.1 Jailbreak Attack Method
We evaluate PSA methods under the attack of the following jailbreak methods.
  • AutoDAN (Liu et al., 2024c): automatically generate stealthy jailbreak prompts by the carefully designed hierarchical genetic algorithm.
  • PAIR (Chao et al., 2024): use an attacker LLM to automatically generate jailbreaks for a separate targeted LLM without human intervention, where the attacker iteratively queries the target LLM to update and refine a candidate jailbreak.
  • GCG (Zou et al., 2023): add a suffix to maximize the probability that the model produces an affirmative response.
  • ICA (Wei et al., 2023b): propose the In-Context Attack (ICA) which employs harmful demonstrations to subvert LLMs.
C.2 Baselines for Post Safety Alignment
We evaluate SafePatching against the following PSA baseline methods.
@@ C.2 Baselines for Post Safety Alignment
  • Gradient Ascent (GA) (Yao et al., 2023): Gradient ascent is performed on harmful dataset during the training process. Furthermore, it also applies a forward KL-divergence with the original backbone on extra general datafor utility keeping.
@@
  • GA + Mismatch (Yao et al., 2023): Built upon GA approach, GA + Mismatch adds random responses from normal dataset for harmful inputs. And another gradient descent optimization is added accordingly to each training step. The forward KL-divergence is also leveraged on extra general data for the preservation of utility.
  • $\text{RESTA}_{d}$ (Bhardwaj et al., 2024): RESTA stands for REstoring Safety through Task Arithmetic. It involves a simple arithmetic addition of a safety vector to the weights of the safety compromised model for post safety alignment. The subscript $d$ refer to the method when the safety vector is added to the SFT model with parameter sparsification method DARE (Yu et al., 2024).
@@
  • Negative Preference Optimization (NPO) (Zhang et al., 2024a): A simple preference-optimization-inspired method that could efficiently and effectively unlearn a target dataset with negative samples. The forward KL-divergence is also applied on extra general data.
@@
  • ROSE (Zhong et al., 2024): A simple PSA method to directly boost the safety of existing aligned LLMs. The principle of it is to improve the probability of desired safe output via suppressing the undesired output induced by the carefully-designed reverse prompts.
@@
  • SafeDecoding (Xu et al., 2024): It is designed to mitigate jailbreak attacks by identifying safety disclaimers and amplifying their token probabilities, while simultaneously attenuating the probabilities of token sequences that are aligned with the objectives of jailbreak attacks. It is important to highlight that, unlike the other decoding-based methods that function solely in the decoding phase without any training, SafeDecoding additionally incorporates a process for training a safety expert.
  • Self-Contrastive Decoding (Self-CD) (Shi et al., 2024): A decoding-based method for over-safety mitigation. It collects the model’s responses to the same question by emphasizing within the prompts whether safety is taken into account. Then it mitigates over-safety by contrasting the output distributions of these answers.
  • SCANS (Cao et al., 2024): It extracts the refusal steering vectors within the activation space and utilizes vocabulary projection to anchor some specific safety-critical layers which influence model refusal behavior. Then it steers the model behavior accordingly.
  • Surgery (Wang et al., 2024): It extracts a false refusal vector and then ablates this vector to reduce false refusal rate.
C.3 Baselines for Model Merging
@@
To further demonstrate the efficacy of our SafePatching in mitigating the conflict between patches for safety and over-safety, while preserving the utility of the backbone, we compare it with current state-of-the-art model merging methods.
@@
  • Average Merging: It combines the parameters of multiple backbone models by averaging them to create a merged model.
@@
  • Task Arithmetic (Ilharco et al., 2022a): It uses a scaling factor to balance the contributions between different models being merged.
@@
  • Fisher Merging (Matena and Raffel, 2022): First, it estimates the significance of the parameters by calculating the Fisher information matrix, and then it merges the parameters according to their importance.
@@
  • TIES-Merging (Yadav et al., 2023): It seeks to resolve parameter conflicts during model merging. Initially, it trims parameters with lower magnitudes and then addresses sign disagreements. Finally, parameters with consistent signs are merged.
BS GA GD GA-LR GD-LR $p$ $a$ $b$ $\alpha$ $\beta$
LLaMA-2-7B-Chat 32 62 steps 141 steps 3.5e-5 5e-5 30 3 2 1 0.2
LLaMA-3-8B-Instruct 32 1 epo 3 epo 5e-6 1e-5 30 3 2 1 0.1
Gemma-1.1-7B-it 32 1 epo 3 epo 2e-5 5e-5 30 3 2 0.01 0.5
Mistral-7B-Instruct-v0.1 32 45 steps 3 epo 1e-5 1e-5 30 3 2 1 0.15
Table 5: Detailed hyper-parameter settings for SafePatching on different backbones. LR is learning rate. BS is batch size. GA and GD are gradient ascent and descent, respectively.
Appendix D Implementation Details
@@
Our experiments are implemented with PyTorch (Paszke et al., 2019) and Transformer library (Wolf et al., 2020). The training process for all backbones is performed on 4 NVIDIA Tesla A100 using DeepSpeed (Rasley et al., 2020) repository with ZeRo-2 optimization. And the evaluation for utility is performed with lm-evaluation-harness (Gao et al., 2023) and GPT-4o for MT-Bench.
Experimental Setting for Attack Setup
For GCG (Zou et al., 2023), AutoDAN (Liu et al., 2024c), and PAIR (Chao et al., 2024), we follow the approach outlined in (Chao et al., 2024; Zeng et al., 2024) and use 150 distinct vanilla harmful queries from Advbench (Zou et al., 2023) to craft specific attack prompts for each model. Of these, 100 are selected for the Seen setting and 50 for the Unseen setting. The hyper-parameters are set according to the original papers. For ICA (Wei et al., 2023b), we directly use the pre-configured template prompt with 100 data points for the Seen setting and 50 data points for the Unseen setting.
Experimental Setting for Post Safety Alignment
Here the harmful training data $D_{h}$ is formed by those harmful questions that are successfully jailbroken by four jailbreak methods. This is a simulation of SafePatching for “post-hoc” remediation of the jailbreak attack. Specifically, we collect 500 jailbroken data for training in total. We strictly control the training steps of gradient ascent to prevent the resulting model from generating “hallucination-style” responses such as whitespace due to excessive optimization. And all backbones are trained with their official chat template. We do not adopt their safety system prompt because we need to eliminate this factor to prove that the improvement in safety performance is entirely due to our SafePatching. Detailed hyper-parameter settings for SafePatching on different backbone models are shown in Table 5.
@@
Further, we carefully evaluate the official implementations of all baselines, in order to make the comparison as fair as possible. We strictly follow the hyper-parameter settings in their original code, where maximum training step of gradient ascent is set to 200. Following (Yao et al., 2023), TruthfulQA (Lin et al., 2022) is adopted for the general training data $D_{g}$. In $\text{RESTA}_{d}$, this is replaced with English Alpaca (Taori et al., 2023) to align with their original settings. If this could not reach the expected performance, we carry out the hyper-parameter search of the learning rate and batch size. We also perform hyper-parameter adjustment for all model merging baselines, and the optimal results are reported in Table 2.
Method Accuracy
Longformer-based 96.70%
GPT-4o-based 97.80%
Table 6: Comparison of Longformer-based and GPT-4o-based methods in terms of Accuracy.
Appendix E Automatic Evaluation of Harmful Rate
As pointed out by Wang et al. (2023), a key limitation of GPT-4-based evaluation is data privacy, as the model cannot be deployed locally. To address this problem, they introduce PLM-based evaluators. To be more specific, a Longformer-based (Beltagy et al., 2020) classifier is fine-tuned by Wang et al. (2023) over human annotations for each input-output harmful data to perform harmful response detection. Each response to be evaluated will be classified into six categories as shown in Table 4, with only Label 5 being considered as a harmful response. In our work, we directly adopt it to calculate the ASR in our experiments.
To validate the accuracy and reliability of the Longformer-based classifier used in our experiments, we sample 50 instances from AdvBench and compare its classification performance with GPT-4o, where human judgements serve as the golden reference. Results are shown in Table  6. This comparison confirms the Longformer-based model’s effectiveness in our specific setting.
Appendix F Over-Safety Evaluation
XSTest
comprises 250 safe prompts across ten different prompt types that well-calibrated models should not refuse to comply with.
OKTest
contains 300 testing samples that are characterized by safe questions containing harmful and sensitive words.
Evaluation Protocol
@@ Appendix F Over-Safety Evaluation
Following Röttger et al. (2024), the LLM backbone would generate three main types of responses on the XSTest and OKTest: (1) Full Compliance, which is the preferred outcome for safe prompts. This includes any direct attempt to answer the question, regardless of accuracy or relevance. (2) Full Refusal, which is an undesirable outcome for safe prompts. In this case, the model explicitly states that it will not answer the question and may offer alternative suggestions instead. (3) Partial Refusal, which combines elements of both refusal and compliance. This includes responses that initially refuse to answer but then go on to address the question. And we only count responses fall into the Full Refusal as the refusal rate to measure the performance of over-safety. Besides, the participants in our human evaluation are volunteered transparently informed of our research intent, with reasonable wages paid.
@@
AutoDAN GCG ICA PAIR AutoDAN GCG ICA PAIR XSTest OKTest AVG. MT-Bench
LLaMA-2-7B-Chat 19.00 46.00 4.00 27.00 12.00 44.00 18.00 10.00 8.00 4.67 40.35 6.01
GA 12.00 20.00 2.00 15.00 10.00 20.00 6.00 4.00 35.20 38.67 39.02 4.01
GA + Mismatch 20.00 45.00 4.00 20 16 46.00 12.00 6.00 22.40 25.33 38.42 4.56
$\text{RESTA}_{d}$ 58.00 42.00 3.00 56.00 62.00 28.00 4.00 48.00 66.33 41.60 39.39 5.49
NPO 27.00 36.00 2.00 45.00 18.00 40.00 40.00 6.00 18.80 18.67 39.26 5.62
SafeDecoding 2.00 2.00 0 4.00 2.00 0 0 4.00 80.80 59.67 - 5.72
ROSE 1.00 0 0 3.00 4.00 0 0 0 43.20 40.33 - 4.14
Self-CD 2.00 40.00 0 8.00 6.00 40.00 4.00 0 11.60 15.60 - 3.98
SCANS 23.00 52.00 3.00 29.00 66.00 22.00 2.00 30.00 33.60 5.67 - 5.84
Surgery 27.00 60.00 1.00 35.00 28.00 60.00 0 30.00 11.60 4.33 - 5.45
SafePatching (Ours) 10.00 8.00 2.00 9.00 10.00 12.00 2.00 6.00 7.60 2.33 40.42 6.14
Table 7: The overall results on the safety, over-safety and utility benchmarks with LLaMA-2-7B-Chat backbone. The evaluation metrics for safety and over-safety are ASR and refusal rate, respectively. We report the average results on all the utility datasets except for MT-bench (1 - 10).
[…]
AutoDAN GCG ICA PAIR AutoDAN GCG ICA PAIR XSTest OKTest AVG. MT-Bench
LLaMA-3-8B-Instruct 28.00 4.00 14.00 4.00 38.00 12.00 24.00 0 2.80 9.67 59.31 8.20
GA 12.00 20.00 2.00 15.00 10.00 20.00 6.00 4.00 35.20 38.67 39.02 4.01
GA + Mismatch 20.00 45.00 4.00 20.00 16.00 46.00 12.00 6.00 22.40 25.33 38.42 4.56
$\text{RESTA}_{d}$ 32.00 4.00 0 10.00 62.00 28.00 4.00 48.00 66.33 41.60 39.39 5.49
NPO 27.00 36.00 2.00 45.00 18.00 40 40.00 6.00 18.80 18.67 39.26 5.62
ROSE 0 0 0 2.00 0 0 2.00 0 43.20 40.33 - 5.74
Self-CD 0 6.00 0 10.00 0 0 8.00 0 10 18.67 - 5.06
SCANS 92.00 0 59.00 8.00 90.00 0 66.00 16.00 9.60 12.67 - 7.92
Surgery 9.00 44.00 0 8.00 0 2.00 0 20.00 7.60 4.00 - 6.33
SafePatching (Ours) 15.00 2.00 0 2.00 28.00 12.00 14.00 0 1.60 6.33 59.39 8.18
Table 8: The overall results on the safety, over-safety and utility benchmarks with LLaMA-3-8B-Instruct backbone. The evaluation metrics for safety and over-safety are ASR and refusal rate, respectively. We report the average results on all the utility datasets except for MT-bench (1 - 10).
[…]
AutoDAN GCG ICA PAIR AutoDAN GCG ICA PAIR XSTest OKTest AVG. MT-Bench
Gemma-1.1-7B-it 39.00 20.00 29.00 10.00 50.00 28.00 46.00 12.00 28.80 23.33 51.16 6.82
GA 13.00 0 14.00 10.00 26.00 22.00 50.00 0 30.40 38.67 46.41 3.96
GA + Mismatch 18.00 2.00 12.00 10.00 30.00 26.00 54.00 0 32.40 40.67 45.74 4.23
$\text{RESTA}_{d}$ 32.00 34.00 8.00 14.00 20.00 44.00 12.00 12.00 22.33 27.20 47.01 5.71
SCANS 72.00 24.00 12.00 10.00 70.00 22.00 12.00 16.00 5.20 8.67 - 6.86
Surgery 2.00 0 0 3.00 2.00 0 0 8.00 63.20 70.00 - 5.56
SafePatching (Ours) 0 2.00 0 5.00 0 20.00 40.00 0 16.00 14.67 51.27 6.94
Table 9: The overall results on the safety, over-safety and utility benchmarks with Gemma-1.1-7B-it backbone. The evaluation metrics for safety and over-safety are ASR and refusal rate, respectively. We report the average results on all the utility datasets except for MT-bench (1 - 10).
[…]
AutoDAN GCG ICA PAIR AutoDAN GCG ICA PAIR XSTest OKTest AVG. MT-Bench
Mistral-7B-Instruct-v0.1 97.00 38.00 90.00 75.00 92.00 34.00 64.00 88.00 14.00 6.33 45.67 6.49
GA 8.00 0 0 0 56.00 2.00 10.00 6.00 100 100 43.99 3.12
GA + Mismatch 17.00 0 0 0 60.00 2.00 2.00 4.00 100 100 44.63 3.46
$\text{RESTA}_{d}$ 25.00 54.00 74.00 63.00 28.00 48.00 82.00 68.00 1.67 2.00 41.30 3.38
ROSE 91.00 44.00 83.00 39.00 98.00 42.00 38.00 78.00 14.40 6.33 - 3.89
Self-CD 95.00 82.00 88.00 92.00 96.00 50.00 96.00 84.00 14.00 11.33 - 3.57
SCANS 65.00 60.00 69.00 35.00 66.00 46.00 60.00 30.00 51.20 37.67 - 6.01
Surgery 48.00 2.00 23.00 6.00 50.00 2.00 30.00 10.00 26.00 38.00 - 5.21
SafePatching (Ours) 23.00 0 3.00 1.00 56.00 2.00 2.00 2.00 5.20 3.33 45.29 6.38
Table 10: The overall results on the safety, over-safety and utility benchmarks with Mistral-7B-Instruct-v0.1 backbone. The evaluation metrics for safety and over-safety are ASR and refusal rate, respectively. We report the average results on all the utility datasets except for MT-bench (1 - 10).
@@
Table 11: Training data required by different methods.
@@
Table 12: ATGR of different methods to manifest the inference cost.
Appendix G Additional Experimental Results
G.1 Results on Different Backbones
The experimental results of the current PSA baseline methods based on unlearning and decoding on the LLaMA-2-7B-Chat, LLaMA-3-8B-Instruct, Gemma-1.1-7B-it and Mistral-7B-Instruct-v0.1 are shown in Tables 7, 8, 9 and 10. We do not report the results of SafeDecoding on the LLaMA-3-8B-Instruct, Gemma-1.1-7B-it and Mistral-7B-Instruct-v0.1 backbones because this method is only performed on LLaMA-2-7B-Chat in the original paper and it requires using a portion of safe data to train a safety-enhanced expert model. However, this portion of data is not publicly available, so we are unable to reproduce the results. And results of Self-CD and ROSE are not provided on Gemma-1.1-7B-it because this backbone does not support the customized system prompt.
The results demonstrate that our SafePatching is still capable of achieving the three goals of enhancement of safety, mitigation of over-sensitivity, and maintenance of utility. Similarly, SafePatching continues to improve the model’s multi-turn instruction-following ability on the MT-bench, enhancing the quality of the model’s responses to user inputs by reducing its over-sensitivity to safety concerns. These extensive experimental results demonstrate that SafePatching can serve as an effective and scalable PSA solution. We hope that our method can provide inspiration for future work in achieving comprehensive PSA.
G.2 Comparison of Training and Inference Costs